OpenStudy (anonymous):

Divide Anyone want to help me out? x^2 -4/x-8/x-2 A.(x-2)(x+2)/x -8 B.x-8/x+2 <-- This is wrong C.x-2/x-8 D.x+2/x-8 I think it is C i'm just being for sure I failed math the first time trying not to do that again. XD

Nnesha (nnesha):

is it \[\huge\rm \frac{ \frac{ x^2 -4}{ x-8 } }{ x-2 }\]

OpenStudy (anonymous):

Yes it is.

Nnesha (nnesha):

alright show your work plz

Nnesha (nnesha):

or explain how did you get C ?

OpenStudy (anonymous):

To be honest guess it XD

Nnesha (nnesha):

i knew it don't guess and btw it's too hard to pick the option whn there are 4!

OpenStudy (anonymous):

I Know but if you don't know how to solve it it's the next option.

Nnesha (nnesha):

\[\huge\rm \frac{ \frac{ x^2 -4}{ x-8 } }{ x-2 }\] first of all change division the multiplication so multiply first fraction by the reciprocal of the bottom fraction example \[\large\rm \frac{ \frac{ a }{ b } }{ \frac{ c }{ d }}\] \[\bf \frac{ a }{ b } \times \frac{ d }{ c }\]

OpenStudy (anonymous):

4 Like this